# Tailored Emotional LLM-Supporter: Enhancing Cultural Sensitivity

### v1

@@ Abstract
Large language models (LLMs) show promise in offering emotional support and generating empathetic responses for individuals in distress, but their ability to deliver culturally sensitive support remains underexplored due to lack of resources. In this work, we introduce CultureCare, the first dataset designed for this task, spanning four cultures and including 1729 distress messages, 1523 cultural signals, and 1041 support strategies with fine-grained emotional and cultural annotations. Leveraging CultureCare, we (i) develop and test four adaptation strategies for guiding three state-of-the-art LLMs toward culturally sensitive responses; (ii) conduct comprehensive evaluations using LLM judges, in-culture human annotators, and clinical psychologists; (iii) show that adapted LLMs outperform anonymous online peer responses, and that simple cultural role-play is insufficient for cultural sensitivity; and (iv) explore the application of LLMs in clinical training, where experts highlight their potential in fostering cultural competence in future therapists.111Github: https://github.com/UKPLab/arxiv2025-culturecare.
Content Warning: This paper includes examples that some readers may find offensive or triggering. These instances are presented for research purposes only and do not represent the views of the authors or affiliated institutions.
@@ 1 Introduction
Large language models (LLMs) have shown growing potential in offering emotional support, with recent work demonstrating that LLMs can provide empathetic, contextually relevant responses for individuals experiencing distress Zheng et al. (2024); Zhan et al. (2024); Ye et al. (2025). This emerging capability is particularly promising in online spaces, where peer support communities play a vital role in helping individuals navigate emotional challenges. However, culture shapes human emotional experiences and influences the stressors people encounter in daily life Markus and Kitayama (1991); Mesquita and Frijda (1992); Chun et al. (2006); Mesquita et al. (2017). As a result, one critical dimension remains underexplored — the cultural sensitivity of these LLMs’ support responses.
@@
Effective emotional support is deeply shaped by cultural knowledge, unspoken assumptions, norms, and values that influence how distress is expressed and how support is received Taylor et al. (2007); Matsumoto et al. (2008); Kim et al. (2008). However, when LLMs are used as tools for emotional support, they may struggle to recognize these culturally embedded cues of distress in the first place Aleem et al. (2024). Even well-intentioned responses can cause harm or alienation without cultural grounding and sensitivity, rather than providing the support users seek Lissak et al. (2024); Moore et al. (2025). For example, in a collectivist culture, an LLM advising a user to cut ties with their family in pursuit of personal happiness, without acknowledging the cultural weight of familial duty, could come across as offensive or morally unacceptable. Similarly, in cultures where mental health is heavily stigmatized, a blunt recommendation to “seek therapy immediately” might intensify feelings of shame or social isolation rather than offering relief.
@@
Recent research tries to address this issue through a focused case study on Pakistani culture Aleem et al. (2024). However, the study offers limited generalizability due to its narrow set of manually created distress scenarios, leaving the broader challenge of culturally aware emotional support largely unexplored. The lack of suitable datasets and evaluation of adaptation methods has hindered progress in this area. Therefore, in this work, we address the research questions by presenting a comprehensive multi-cultural investigation into LLMs’ ability to provide culturally sensitive emotional support.
@@
To address this data gap, we introduce CultureCare. To the best of our knowledge, this is the first dataset designed to support the study of culturally-aware peer emotional support (Figure 1). CultureCare spans four distinct global cultures, namely Arabic, Chinese, German, and Jewish, and fine-grained annotated for both emotional support strategies and culturally-relevant signals, as shown in Table 1. The dataset consists of distress-response pairs sourced from real-world support interactions from Reddit, annotated for support type and cultural signals, allowing for nuanced development and evaluations for adapting LLMs’ responses across cultural contexts. Using CultureCare, we evaluate three state-of-the-art LLMs with tailored prompting strategies for adaptation. Through both automated evaluations and in-culture human evaluations, we find that while incorporating basic cultural information helps, a more effective adaptation requires detailed guidelines and attention to contextualized, explicit cultural signals. These LLM-generated responses are often better than human responses online.
@@
While our primary focus is on peer support — where the LLM acts as a supportive peer rather than providing professional help — we also explore its potential in clinical training settings, namely in training psychology students for culturally sensitive therapy. Expert feedback highlights strong safety and promising utility.
@@
To sum up, our contributions are: 1. We release the first dataset — CultureCare — for evaluating and training LLMs in culturally-aware emotional support, spanning four cultures with fine-grained annotations. The dataset comprises 1729 annotated distress messages, 1523 cultural signals, and 1041 support strategies. 2. We develop and test four adaptation strategies to guide three popular state-of-the-art LLMs toward generating culturally-sensitive support responses. 3. We provide comprehensive evaluations involving a LLM judge, in-culture human evaluators, and clinical psychologists to assess both emotional and cultural quality of generated responses.
2 Related Work
Culture and mental health. Culturally sensitive counselling is a well-established consideration in clinical psychology and healthcare settings (Bernal et al., 1995; Resnicow et al., 1999; Kreuter and McClure, 2004; Taylor et al., 2007; Tao et al., 2015, among others). Prior research has explored various aspects of incorporating cultural sensitivity in practical domains outside of NLP, including the importance of cultural humility in improving therapy outcomes Owen et al. (2016), disparities in engagement and follow-up care across demographics Zeber et al. (2017), and the need to embed cultural competence in training programs Benuto et al. (2018). However, the application of LLMs in culturally sensitive mental health remains limited. Focusing on formal therapy settings, recent work Abbasi et al. (2025); Kim et al. (2025) explores LLM-generated synthetic multi-turn clinical conversations in multilingual contexts. Their focus on clinical therapy and synthetic data generation differs from ours, which centers on peer emotional support and adaptation strategies.
LLMs for emotional support. Existing work has shown that LLMs can provide empathetic and supportive responses when appropriately guided Zhan et al. (2024). Studies such as Liu et al. (2021); Zheng et al. (2024); Zhang et al. (2024) examine how LLMs respond to distress messages using various support strategies. However, they largely overlook the influence of cultural context in shaping emotional needs and support preferences. While Aleem et al. (2024) considers cultural context, its focus on a small set of scenarios from Pakistani culture limits its generalizability.
Furthermore, recent studies on LLMs in cognitive behavioural therapy Goel et al. (2025); Zhou et al. (2025); Zhang et al. (2025) emphasize clinical settings with structured treatment frameworks, offering little attention to cultural nuance. In contrast, our work focuses on peer-to-peer emotional support, where cultural understanding is essential for effective and supportive communication. We address this gap by incorporating cultural signals from diverse cultural communities into our annotation process, and by assessing emotional support responses from humans and LLMs, based not only on empathy but also on their cultural relevance.
3 CultureCare
We present CultureCare, a multi-cultural dataset with fine-grained span-level annotation of cultural signals and emotional distress.222We deliberately chose not to use language as the defining boundary of culture, recognizing that culturally influenced distress can be expressed in any language in online communities. As a result, CultureCare includes posts both in English and in the native languages associated with each culture. The language distribution varies by culture; see Appendix A.5. An overview of CultureCare’s construction pipeline is in Figure 2. Briefly, we begin by collecting culturally relevant mental health data from Reddit using targeted search queries. We then apply a combination of rule-based and LLM-based filters to remove noise. Finally, in-culture annotators and reviewers select the spans and label the data for emotional distress, cultural signals, and support strategies. The annotation categories for each dimension are detailed in Appendix A.4.
Table 1 compares CultureCare with existing resources. CultureCare is distinctive in its focus on multiple cultures, the explicit annotation of cultural signals and emotional distress.
@@ 3.1 Data Collection and Filtering
Candidate posts. We use Reddit as our dataset source due to its global user base and peer-driven mental health discussions. Our data is collected via the Reddit API333https://www.reddit.com/dev/api/
Data collection and annotations are approved by the ethics board.. We focus on two setups: 1. searching mental health subreddits with culture-specific keywords, and 2. searching culture-specific subreddits with mental health keywords. The list of subreddits and keywords is in Appendix A.2. For each relevant post, we fetch the top-voted comment as the ideal supportive response, as it typically offers emotional support and resonates with readers, reflected in its high upvotes.
Filtering. Our initial dataset contained 9160 posts, many of which were noisy — e.g., general mental health tips, reactions to global events, or content unrelated to the target cultures. Since manual review of all posts was infeasible, we first applied rule-based filters to remove both explicit noise (e.g., URLs-only posts) and LLM-based filtering to remove implicit noise (e.g., posts lacking personal distress). This leaves us with 2671 posts for manual review. The full set of filters is detailed in Appendix A.3. After the final manual review, we retained 462 high-quality posts that clearly express culturally-situated emotional distress, with optional responses by humans. While small, this dataset allows for focused, fine-grained annotation.
@@ 3.2 Annotations and Quality Assurance
Each post was annotated by a in-culture annotator using detailed guidelines (Appendix A.4) identifying spans with emotional distress, cultural signals, and emotional support. On average, each post took approximately 10 minutes to annotate, as they were often lengthy and required attentive reading. To ensure high quality, a second in-culture annotator reviewed the initial annotations and provided comments in cases of disagreement; however, such instances were rare and typically involved additions rather than corrections.
@@ 3.3 Analysis of CultureCare
Statistics. CultureCare includes 462 posts (90% with responses444The rest 10% had deleted comments or no comments.), containing 1729 annotated distress messages, 1523 cultural signals, and 1041 support strategies. We also extract demographic details (e.g., age, gender, religion) from the posts using an LLM, yielding 926 signals. Example annotated posts are in Appendix F.
Prevalent norms and values. Appendix A.7 illustrates how culture shapes emotional distress, highlighting recurring themes such as family control, mental health invalidation, and gender expectations. Some experiences are culture-specific, including LGBTQ+ rejection in Arabic contexts and financial stress in German ones. While emotional struggles are universal, their expression and root causes depend on the culture.
Cultural signals types and support strategies. Arabic posts emphasize values and explicitly state demographic details, while Chinese and German prioritize norms and morals. Jewish posts focus heavily on concepts, like rabbi. Artifacts, knowledge, and language have a minimal presence in all cultures. In human responses, providing suggestions is most common across all cultures. We also observe that while Arabic and Chinese responses emphasize self-disclosure and affirmation, German responses are more balanced, and Jewish responses more often provide information. Asking questions is the least common strategy.
Demographic diversity. We extract detailed demographic information, revealing various patterns between age, gender, and other attributes.555In some cases, only high-level cultural indicators, such as “Arabic” or “Chinese”, are extractable. For example, extractable Jewish posts are predominantly female (80%), while extractable Arabic and Chinese posts have a high proportion of young adults (47% and 50%, respectively). The extractions also reflect diversity in birthplace, residence, marital status, education level, and profession. These variations highlight the rich cultural and social heterogeneity of the dataset. More details are provided in Appendix A.6.
Intensity and empathy scores. We additionally provide analysis on cultural differences in distress intensity and response empathy. In our data, in general, Jewish posts show the highest intensity, followed by German and Arabic, while Arabic responses demonstrate the highest empathy. See details in Appendix A.9.
@@ 4 Adaptation Methods
We experiment with basic adaptation strategies using prompts and a compound strategy by combining all strategies, prompts in Appendix C.
Standard (no suffix). By default, we prompt the model to be a Redditor, matching the context of the data. This variation serves as a baseline for comparing adaptation strategies.
Culture-informed role-playing (+culture). Building on prior research AlKhamissi et al. (2024); Tao et al. (2024), instructing LLM to role-play the cultural background of the person is a simple yet effective method for aligning LLM responses with culturally relevant values. Hence, this could enable more empathetic, appropriate responses, removing the cultural difference barrier in empathy Cikara et al. (2014); Davis (2018).
@@
Explicit cultural signals (+annotation). We explicitly add the data annotations of posts from our dataset to the prompt. The goal of this strategy is to understand whether explicitly providing LLMs with richer contextual information can improve understanding and output in an implicit cross-cultural setting.
Combined (+cga). In this method, we combined the above three basic strategies, culture, guided, and annotation (Figure 11, Appendix C). We modified the guidelines in the +guided strategy by removing CCCI-R items that focus on cross-cultural differences. Here, the LLM will be provided with explicit information and guidelines, as well as role-playing a person from the same culture.

### v2

@@ Abstract
Large language models (LLMs) show promise in offering emotional support and generating empathetic responses for individuals in distress, but their ability to deliver culturally sensitive support remains underexplored due to a lack of resources. In this work, we introduce CultureCare, the first dataset designed for this task, spanning four cultures and including 1729 distress messages, 1523 cultural signals, and 1041 support strategies with fine-grained emotional and cultural annotations. Leveraging CultureCare, we (i) develop and test four adaptation strategies for guiding three state-of-the-art LLMs toward culturally sensitive responses; (ii) conduct comprehensive evaluations using LLM-as-a-Judge, in-culture human annotators, and clinical psychologists; (iii) show that adapted LLMs outperform anonymous online peer responses, and that simple cultural role-play is insufficient for cultural sensitivity; and (iv) explore the application of LLMs in clinical training, where experts highlight their potential in fostering cultural competence in novice therapists. Github: https://github.com/UKPLab/eacl2026-culturecare.111Content Warning: This paper includes examples that some readers may find offensive or triggering. These instances are presented for research purposes only and do not represent the views of the authors or affiliated institutions.
@@ 1 Introduction
Large language models (LLMs) have shown growing potential in offering emotional support, with recent work demonstrating that LLMs can provide empathetic, contextually relevant responses for individuals experiencing distress Zheng et al. (2024); Zhan et al. (2024); Ye et al. (2025). This emerging capability is particularly promising in online spaces, where peer support communities play a vital role in helping individuals navigate emotional challenges. However, culture shapes human emotional experiences and influences the stressors people encounter in daily life Markus and Kitayama (1991); Mesquita and Frijda (1992); Chun and others (2006); Mesquita et al. (2017). As a result, one critical dimension remains underexplored—the cultural sensitivity of these LLMs’ support responses.
@@
Effective emotional support is deeply shaped by cultural knowledge, unspoken assumptions, norms, and values that influence how distress is expressed and how support is received Taylor et al. (2007); Matsumoto et al. (2008); Kim et al. (2008). However, when LLMs are used as tools for emotional support, they may struggle to recognize these culturally embedded cues of distress in the first place Aleem et al. (2024). Even well-intentioned responses can cause harm or alienation without cultural grounding and sensitivity, rather than providing the support users seek Lissak et al. (2024); Moore et al. (2025). For example, in a collectivist culture, an LLM advising a user to cut ties with their family for personal happiness, without acknowledging the cultural weight of familial duty, may seem offensive or immoral. Similarly, in cultures where mental health is heavily stigmatized, bluntly urging someone to “seek therapy” might intensify feelings of shame or social isolation rather than offering relief.
@@
Recent research tries to address this issue through a case study on Pakistani culture Aleem et al. (2024). However, the study offers limited generalizability due to its narrow set of manually created distress scenarios, leaving the broader challenge of culturally sensitive emotional support largely unexplored. The lack of suitable datasets and the evaluation of adaptation methods have hindered progress in this area. Therefore, here, we present a comprehensive multi-cultural investigation into LLMs’ ability to provide culturally sensitive emotional support.
@@
To address the data gap, we introduce CultureCare. To the best of our knowledge, this is the first dataset designed to support the study of culturally-sensitive peer emotional support (Figure 1). CultureCare spans four distinct global cultures, namely Arabic, Chinese, German, and Jewish, and collected fine-grained annotations for both emotional support strategies and culturally-relevant signals, as shown in Table 1. The dataset consists of distress-response pairs sourced from real-world interactions from Reddit, annotated for support type and cultural signals, allowing for nuanced development and evaluations for adapting LLMs’ responses across cultural contexts. Using CultureCare, we evaluate three state-of-the-art LLMs with tailored prompting strategies for adaptation. Through both automated evaluations and human evaluations, we find that while incorporating basic cultural information helps, a more effective adaptation requires detailed guidelines and attention to contextualized, explicit cultural signals.
@@
While our primary focus is on peer support—where the LLM acts as a supportive peer rather than providing professional help—we also explore its potential in clinical training settings of training psychology students to conduct culturally competent therapy Benuto et al. (2018). Expert feedback highlights strong safety and promising utility.
@@
To sum up, our contributions are: First, we release the first dataset, CultureCare, for evaluating and adapting LLMs in culturally-sensitive emotional support, spanning four cultures with fine-grained annotations. The dataset comprises 1729 annotated distress messages, 1523 cultural signals, and 1041 support strategies; second, we develop and test four adaptation strategies to guide three popular state-of-the-art LLMs toward generating culturally-sensitive support responses; and third, we provide comprehensive evaluations involving LLM-as-a-Judge, in-culture human evaluators, and clinical psychologists to assess both the emotional and cultural aspects of the generated responses.
@@
2 Related Work
Culture and mental health. Culturally sensitive counselling is a well-established consideration in clinical psychology and healthcare settings (Bernal and others, 1995; Resnicow et al., 1999; Kreuter and McClure, 2004; Taylor et al., 2007; Tao et al., 2015, among others). Prior research has explored various aspects of incorporating cultural sensitivity in practical domains outside of NLP, including the importance of cultural humility in improving therapy outcomes Owen et al. (2016), disparities in engagement and follow-up care across demographics Zeber et al. (2017), and the need to embed cultural competence in training programs Benuto et al. (2018). However, the application of LLMs in culturally sensitive mental health remains limited. Focusing on formal therapy settings, recent work Abbasi et al. (2025); Kim et al. (2025) explores LLM-generated synthetic clinical conversations in multilingual contexts. Their focus on clinical therapy and synthetic data generation differs from ours, which centers on examining LLMs for culturally-aware emotional support across several LLMs and adaptation strategies.
LLMs for emotional support. Existing work has shown that LLMs can provide empathetic and supportive responses when appropriately guided Zhan et al. (2024). Studies such as Liu et al. (2021); Zheng et al. (2024); Zhang et al. (2024) examine how LLMs respond to distress messages using various support strategies. However, they largely overlook the influence of cultural context in shaping emotional needs and support preferences. While Aleem et al. (2024) considers cultural context, its focus on a small set of scenarios from the Pakistani culture limits its generalizability. Furthermore, recent studies on LLMs in cognitive behavioural therapy Goel et al. (2025); Zhou et al. (2025); Zhang et al. (2025) emphasize clinical settings with structured treatment frameworks, offering little attention to cultural nuance. In contrast, our work focuses on emotional support where cultural understanding is essential for effective and supportive communication. We address this gap by incorporating cultural signals from diverse cultural communities into our annotation process, and by assessing emotional support responses from both humans (Reddit users) and LLMs, based not only on evaluation metrics like empathy, but also on cultural metrics like showing an understanding of the cultural context.
3 CultureCare
We present CultureCare, a multi-cultural dataset with fine-grained span-level annotation of cultural signals and emotional distress.222We deliberately chose not to use language as the defining boundary of culture, recognizing that culturally influenced distress can be expressed in any language in online communities. As a result, CultureCare includes post-responses pairs both in English and in the native languages associated with each culture. The language distribution varies by culture; see Appendix A.5. Briefly, we begin by collecting publicly available Reddit333An anonymous platform where user identities remain hidden. posts at the intersection of culture and mental health. Specifically, we draw from mental health subreddits using culture-related keywords (e.g., r/depression with the keyword “Chinese”) and vice versa (e.g., r/china with the keyword “depression”). We then apply a combination of rule-based and LLM-based filters to remove noise (i.e., irrelevant posts for the task we target). Every instance consists of a post-response pair. Finally, in-culture annotators annotate these instances along the following dimensions: emotional distress and their intensity, cultural signals and their categories, support messages and their strategies, and overall empathy level. This is followed by quality checks by in-culture reviewers. To protect annotators’ mental health, we provide clear content warnings for sensitive topics and allow annotators to stop the task at any time. The construction pipeline is illustrated in Figure 2. This research was approved by the ethics committee of the Technical University of Darmstadt (EK 121/2024).
@@ 3.1 Data Collection and Filtering
Candidate posts. We use Reddit as our dataset source due to its global user base and peer-driven mental health discussions. Our data is collected via the Reddit API444https://www.reddit.com/dev/api/. We focus on two setups: 1. searching mental health subreddits with culture-specific keywords, and 2. searching culture-specific subreddits with mental health keywords. The list of subreddits and keywords is in Appendix A.2. For each relevant post, we fetch the top-voted comment as the ideal supportive response, as it typically offers emotional support and resonates with readers, reflected in its high upvotes.
Filtering. Our initial dataset contained 9160 posts, many of which were noisy—e.g., general mental health tips, reactions to global events, or content unrelated to the target cultures. Since manual review of all posts was infeasible, we first applied rule-based filters to remove both explicit noise (e.g., URLs-only posts) and LLM-based filtering to remove implicit noise (e.g., posts lacking personal distress). This leaves us with 2671 posts for manual review. The full set of filters is detailed in Appendix A.3. After the final manual review, where annotators were asked to flag any remaining irrelevant posts and not annotate it (e.g., a post from someone who lives in Germany on one of the German subreddits, but is not culturally German), we retained 462 high-quality instances. While small, this dataset allows for focused, fine-grained annotation, and is the largest of its kind (more discussions in Appendix A.10).
@@ 3.2 Annotations and Quality Assurance
Each post-response was annotated by an in-culture annotator, who selects the spans and labels the data for the following dimensions: emotional distress (to highlight the span of text where the poster expresses emotional distress), intensity of emotional distress (how intense are the emotions in the post), cultural signals (to highlight any text indicative of a culture reference in both the post and the response), types of cultural signals (to categorize the annotated cultural signal, e.g., values), support messages (to highlight the span of text in the response where the responder offers support), the support strategy behind the message (to categorize the annotated support message, e.g., offering suggestions) and its empathy level (how empathetic is the support message in the response). To ensure high-quality annotations, an in-culture quality checker reviewed the initial annotations and left comments when they disagreed; the annotator then reviewed and approved these comments. Most additions were suggestions for extra annotations rather than indications of inconsistencies. On average, each post took approximately 10 minutes to annotate and 5-7 minutes to review, as they were often lengthy and required attentive reading.
The definitions and categories for these dimensions are detailed in Appendix A.4, along with guidelines and measures for content sensitivity.
@@ 3.3 Analysis of CultureCare
Statistics. CultureCare includes 462 posts55590% with responses. The remaining 10% had deleted comments or no comments., containing 1729 annotated distress messages, 1523 cultural signals, and 1041 support strategies. We also extract demographic details (e.g., age, gender, religion) from the posts using an LLM, yielding 926 additional annotations. An overview of these statistics per culture is in Table 2, and sample annotated posts are in Appendix F.
Prevalent cultural signals and support strategies by culture. In order to understand the most frequent culture category and emotional support strategy, we compute the number of occurrences of the category, i.e., type, of every culture signal (namely, concepts, knowledge, values, norms and morals, language slang, artifacts, and demographics; definitions and examples in Table 7), and every support message strategy (questions, restatement, reflection of feelings, self-disclosure, affirmation, suggestions, information; definitions and examples in Table 8), annotated by in-culture annotators and approved by quality checkers. We found that the most recurring cultural signals are as follows: Arabic (values), Chinese (norms and morals), German (norms and morals), and Jewish (concepts). Moreover, the most recurring support strategy is “providing suggestions” for all cultures. The full distributions of these categories are in Figure 7 in Appendix A.8.
Demographic diversity. Using GPT-4o-mini, we extract (when present) detailed demographic information 666This goes beyond the culture-related demographic information in our human annotations., namely place of residence, gender, age, born in, marital status, number of people in the household, education, profession, employment, class, and religion. Every extraction comes with both an answer and evidence. For example, out of the evidence (span of text) “I’m a 17M living in Sudan”, the model extracts gender:male, age:17. When any of the demographic fields is not found, the LLM returns “unknown” for that field. We manually inspected a subset of the extractions and found no evidence of inconsistencies or hallucinations. For 50% of the dataset, demographic information could be extracted. We release these LLM-derived annotations alongside the dataset. Finally, we provide a qualitative analysis of the demographic diversity in CultureCare across cultures. The dataset encompasses a broad range of cultural backgrounds and life circumstances, including diverse geographic origins (e.g., Arabic: from Syria, Egypt, Saudi Arabia, and more), age groups (e.g., Chinese: ranging from 15 to 58 years), and professions (e.g., German: working as pizza couriers, social workers, opticians, etc.). This reflects substantial cultural and social heterogeneity within each group. Additional details are in Table 12, Appendix A.6, and the released dataset.
Prevalent norms, morals, and values. In this analysis, we focus on the cultural signals human-annotated in the posts and responses that were categorized, by the in-culture annotators, specifically under “norms and morals” and “values”. To understand culture-specific themes under this category, we prompt GPT-4o, to cluster, per culture, the spans annotated as norms and values, and return the top 5 clusters (by descending order of size). We manually inspect the resulting clusters and find no inconsistencies. The top clusters per culture are shown, with examples for every cluster, in Table 3. This analysis shows that emotional struggles are universal, but their expression and underlying causes vary across cultures, revealing both shared and distinct themes.
| Culture | Themes |
| LGBTQ+ Rejection: came out … as gay … I was not met with acceptance | |
| Gender Roles: …expects a woman to depend fully on a man and her family | |
| Strict Parenting: my dad is extremely strict | |
| Chinese | Mental Health Invalidation: Depression doesn’t exist |
| Family Dynamics: It feels like she only lives her life for me | |
| Strict Parenting: I live with my tiger parents | |
| Financial Stress: I’m on social welfare and feel ashamed | |
| Gender Roles: …men have more responsibilities … may not be … much fun | |
Intensity and empathy scores. Based on the human-annotated intensity level of distress messages in the posts (1: light, 2: moderate, 3: high; definitions in Table 6) and empathy level of support messages in the responses (from 1:not empathetic at all to 5:very empathetic; definitions in Table 9), we compute the average intensity and empathy scores across and per cultures. Our findings show that overall, the average intensity level of the distress messages in our dataset is 1.89, per culture: Arabic (1.81), Chinese (1.77), German (1.98), and Jewish (1.89). The average empathy level overall is 2.77, per culture: Arabic (3.27), Chinese (2.15), German (2.73), and Jewish (3.18). Since these labels were assigned by in-culture annotators (e.g., Arabic annotators labeling Arabic data), they may reflect cultural bias; future work should include out-of-culture annotators to broaden perspectives.
@@ 4 Adaptation Methods
We examine three core prompt-based cultural adaptation strategies, namely role-playing, guided principles, and explicit cultural signals, and a combined approach that integrates all of them, alongside a standard Redditor baseline for comparison.
Standard (redditor). By default, we prompt the model to be a Redditor, matching the context of the data. This variation serves as a baseline for comparing adaptation strategies.
Culture-informed role-playing (+culture). Building on prior research AlKhamissi et al. (2024); Tao et al. (2024), instructing LLM to role-play the cultural background of the person is a simple yet effective method for aligning LLM responses with culturally relevant values. Hence, this could enable more empathetic, appropriate responses, removing the cultural difference barrier in empathy Cikara and others (2014); Davis (1996).
@@
Explicit cultural signals (+annotation). We explicitly add the data annotations of posts from our dataset to the prompt. The goal here is to understand whether explicitly providing LLMs with richer contextual information can improve the response in an implicit cross-cultural setting.
Combined (+cga). In this method, we combine the above three basic strategies, culture, guided, and annotation. We modified the guidelines in the +guided strategy by removing CCCI-R items that focus on cross-cultural differences. Here, the LLM will be provided with explicit information and guidelines, as well as role-playing a person from the same culture.
All adaptation prompts are in Appendix C.